What is the average rate, in miles per hour, of a car that travels 20 miles at 40 miles per hour and then 10 miles at 60 miles per hour?
I also need a formula to along with this. Thank you!

Answer by TimothyLamb(4379)

time per leg:
t = d / s
t = 20/40 + 10/60
t = 1/2 + 1/6
t = 6/12 + 2/12
t = 8/12
---
avg speed:
s = d / t
s = (20 + 10)/(8/12)
s = (30)/(8/12)
s = 45 mph
---
NOTE:
avg speed is NOT: (40+60)/2=50
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php